OpenStudy (anonymous):

I don't understand adding or subtracting a number from a positive or negative number. Could you explain?

OpenStudy (anonymous):

Give an example of one that confuses u...

OpenStudy (anonymous):

Ok. -6-5?

OpenStudy (mimi_x3):

well for example + and + = + - and + = - + and - = - is that what you mean LOLs

OpenStudy (anonymous):

Let's go with an analogy, you borrow 6 dollars (or euros, whatever) from a friend, so you owe your friend 6. Now u borrow another 5, so how much do u owe your freind?

OpenStudy (anonymous):

Eleven right?

OpenStudy (anonymous):

Since they are both positive....or whichever..?

OpenStudy (mimi_x3):

i think its -1 lmao

OpenStudy (anonymous):

Sorry, my computer crashed...

OpenStudy (anonymous):

Yes, 11 is right. Now say it was the other way around, you loaned your friend 11.

OpenStudy (anonymous):

These two 11's are different, right?

OpenStudy (anonymous):

.....then....I'd owe them nothing? [O____o]

OpenStudy (anonymous):

:-) Step at a time....

OpenStudy (anonymous):

Yes, because they are owed to different people yes?

OpenStudy (anonymous):

That's one way of looking at it, from "outside"

OpenStudy (anonymous):

But what about from your point of view...?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

we'd be even?

OpenStudy (anonymous):

How so?

OpenStudy (anonymous):

Since I owed my friend $11 and I gave them $11.

OpenStudy (anonymous):

Or...does that not apply?

OpenStudy (anonymous):

Well, u didn't give it, u loaned it....

OpenStudy (anonymous):

Let's put some signs in to clear things up, let's say u got the loan of 11 and we will say u are now +11

OpenStudy (anonymous):

So your friend would be..?

OpenStudy (anonymous):

So now I have $11 right?

OpenStudy (anonymous):

Friend...um...give you a name?

OpenStudy (anonymous):

:-) If u are +11, what is your friend's position..?

OpenStudy (anonymous):

At zero?

OpenStudy (anonymous):

She gave u 11, so u are +11. Then she is...?

OpenStudy (anonymous):

Minus 11

OpenStudy (anonymous):

Right. And, from "outside" the position is "even"

OpenStudy (anonymous):

Ok. So like this? -6k-5k=-11 -1k=-11 divide by -1 both sides k=11 Am I right?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Let's continue our analogy...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

U pay 4 back to your friend out of the 11 u owe. What's the state of play now, for u, for her, from outside?

OpenStudy (anonymous):

Describe in numbers what has happened...

OpenStudy (anonymous):

Pay 4? You mean I pay her back four? Out of the total of 11 that I owe her?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

For her: I owe her $7 Me: I owe her the 7 Outside(?): I owe her 7 dollars/euros more

OpenStudy (anonymous):

Ok, that's the state of play, now describe in numbers what has happened to get there...

OpenStudy (anonymous):

U had 11....

OpenStudy (anonymous):

Then I gave her four.....and I still owe $7? What happens with the 7?

OpenStudy (anonymous):

So 11-4 =7, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And what are those numbers from your friends point of view?

OpenStudy (anonymous):

She gave me 11. I gave her 4. I still owe her 7 more. Am I right NoW?

OpenStudy (anonymous):

That's from your point of view....ie 11-4 =7 What are the numbers from your friends point of view?

OpenStudy (anonymous):

"I" gave her 11. "She" did not pay me back in full. She gave me 4. But she still owes me 7 more.

OpenStudy (anonymous):

-11 + 4 = -7?

OpenStudy (anonymous):

I believe so. Because "I" do not have the 7 that "She" still owes.

OpenStudy (anonymous):

11-4 =7 Your position -11 + 4 = -7 Your friend Add these two equations, can u?

OpenStudy (anonymous):

Add them together? 7+-7? That would be zero. We'd be even.

OpenStudy (anonymous):

Correct, even though a transaction of 4 took place, the position from outside is still even.

OpenStudy (anonymous):

Now we will complicate matters a bit, your friend says she desperately needs to get 5 of her 7 back, can u help. U haven't got it but u have another friend who agrees to lend u 5 so that u can pay back your friend, so u do that. What are the positions now?

OpenStudy (anonymous):

I give my 1st friend the 5. I only owe her 2 more by this time. My 2nd friend, I now owe HER 5. Either way I owe people 7.

OpenStudy (anonymous):

11-4 +5 -5 =7 Your position -11 + 4 +5 = -2 First friend -5 Second friend Agree?

OpenStudy (anonymous):

Yes, I agree.

OpenStudy (anonymous):

And if u add them up, it will still be zero because all u have done is move money about from one place to another...

OpenStudy (anonymous):

Ok, let's go back to what you posted earlier Ok. So like this? -6k-5k=-11 -1k=-11 divide by -1 both sides k=11 Can u see your mistake now?

OpenStudy (anonymous):

.....no. Not really. My apologies.

OpenStudy (anonymous):

What's -6k+5k?

OpenStudy (anonymous):

-1k

OpenStudy (anonymous):

So then what is -6k -5k?

OpenStudy (anonymous):

1k?

OpenStudy (anonymous):

Think of the minuses as making loans...

OpenStudy (anonymous):

So...-11k?

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

Moving on.. -4 - 7 + 6 is?

OpenStudy (anonymous):

5!

OpenStudy (anonymous):

-4-7 is?

OpenStudy (anonymous):

-11

OpenStudy (anonymous):

-11 +6 is?

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

?

OpenStudy (anonymous):

?!?!? Sorry! I got distracted! -5 is my final answer.

OpenStudy (anonymous):

Right, so -4 - 7 + 6 is -5 (not 5).

OpenStudy (anonymous):

-4-7=-11+6=-5

OpenStudy (anonymous):

We seem to be OK with the idea that -4-7 = -11 so now what about -4 - - 7?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

YOU JUST GRADUATED FROM THIS ROOM!

OpenStudy (anonymous):

Thanks! Um...What does that mean?

OpenStudy (anonymous):

He thinks u have got it..:-) Do u agree?

OpenStudy (anonymous):

MIA, I think she might have....

OpenStudy (anonymous):

I do. But I'll have to review it often or take notes again. I forgot about this after a whole summer. We just got back to school tuesday. Thank you guys! Um...or girls! xD

OpenStudy (anonymous):

ur welcome.

OpenStudy (anonymous):

estudier.... so the correct answer to -6k-5k=-11 k=1?